OpenStudy (anonymous):

factoring a sum/different cubic. a^3+125

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

\[a^3+5^3 = (a+5)(a^2-5a+25)\] you use the formula: \[x^3+y^3 = (x+y)(x^2-xy+y^2)\] where x = a, and y = 5

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

you great

OpenStudy (anonymous):

btw it is best to memorize \[a^3+b^3=(a+b)(a^2-ab+b^2)\] and \[a^3-b^3=(a-b)(a^2+ab+b^2)\] because it is unlikely that you will be able to derive these for yourself on the fly

OpenStudy (anonymous):

and yes. joemath is great!

OpenStudy (anonymous):

ok I will remember. Thanks

OpenStudy (anonymous):

you great too